Explain inductive effect with suitable example

Grade:11

2 Answers

Ravi Agrawal
11 years ago

Partial Transference of Electrons is called Inductive Effect.

It usually occurs due to difference in Electronegativity of bonded atoms.

For example:

  • In CH3Cl, Cl is more electonegative.
  • So, Cl will attract electrons towards it.
  • Here, -I effect is working.
  • -I effect increases the acidity while +I effect decreses the acidity.
  • Simultaneously, -I effect decreases the basicity while +I effect increases the basicity.
Rohan Thomas
7 years ago

The inductive effect is the effect on electron density in one portion of a molecule caused by electron-withdrawing or electron-donating groups elsewhere in the molecule.

In a covalent bond between two atoms of unequal electronegativity, the more electronegative atom draws electron density towards itself. This causes the δ⁺ and δ⁻ charges of the bond dipole.

Hδ+Clδ

If the electronegative atom is joined to a chain of atoms, the positive charge is passed along the other atoms in the chain. This is the electron-withdrawing inductive effect.

The induced change in polarity is less than the original polarity, so the inductive effect rapidly dies out. It is significant only over a short distance.

In NMR spectroscopy, the inductive effect is a factor affecting proton chemical shifts.

If the electron density about a proton is low, the induced field due to electron motions will be weaker than if the electron density is high.

The shielding effect in will be smaller, and a lower external field will be needed for the rf energy to excite the nuclear spin.

The proton is deshielded, and its signal is shifted downfield.

For example, the CH₂ group in propane appears at δ 1.4.

An OH group attached to the CH₂ shifts the signal downfield to δ 3.6.

The effect decreases with distance from the OH group. For example, in butan-1-ol, the CH₂ signals are

CH₃—CH₂—CH₂—CH₂—OH
C-3 = δ1.4; C-2 = δ1.5; C-1 = δ3.6